# Unitarity, Causality, and Solar System Bounds, Significantly Limit Using Gravitational Waves to Test General Relativity

### v3

@@ 3.1.2 Cosmological Constant Counter Term
Let us first consider the cosmological constant, which appears in the action as $S\supset-\int d^{4}x\sqrt{-g}\,\Lambda_{0}$, where we have written the cosmological constant now as $\Lambda\to\Lambda_{0}$ to indicate that this is the bare value. Then consider the following expansion of the determinant of the metric
@@
$=\frac{i}{2}\left(\frac{-i\kappa}{2}\right)\mu^{4-d}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l^{\mu}l^{\nu}-\eta^{\mu\nu}(l^{2}-m^{2})}{l^{2}-m^{2}}$
@@
Then using the Feynman rule that the linear term in the action gives $-i\kappa\Lambda\eta^{\mu\nu}/2$, equation (34) therefore becomes
$-\frac{i\kappa\Lambda_{R}}{2}\eta^{\mu\nu}=-\frac{i\kappa\Lambda_{0}}{2}\eta^{\mu\nu}+\frac{i\kappa m^{4}\eta^{\mu\nu}}{128\pi^{2}}\left(\frac{1}{\epsilon}+\log(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}})+\frac{3}{2}\right).$ (36)
@@ 3.2 Corrections to the Newtonian Potential
$V(\mathbf{r})=\frac{1}{2M_{1}}\frac{1}{2M_{2}}\int\frac{d^{3}q}{(2\pi)^{3}}\,e^{i\mathbf{q}\cdot\mathbf{r}}\mathcal{M}(\mathbf{q})$ (48)
@@
[…]
@@
Figure 1: Correction to the potential $\Delta V$ between point masses due to a scalar ($N=1$) running in a loop versus distance $r$. The black curve is the full numerical result from Fourier transforming Eq. (46). The dashed blue curve is the leading term of the small $m\,r$ result given in Eq. (52). The dashed orange curve is the large $m\,r$ result given in Eq. (56).
@@ 3.2.1 Short Distance Regime
If we take the massless limit we obtain just the single term $\Delta V=-NG^{2}M_{1}M_{2}/(20\pi r^{2})$; this is given as the dashed blue curve in Figure 1. This confirms Ref. Burns:2014bva , but is in disagreement with Ref. Faller:2007sy by a factor of $1/4$. We note that in Ref. Faller:2007sy there is no symmetry factor of $1/2$ in their bubble diagram, and they did not take into account our tadpole Feynman diagrams and the associated renormalization of the cosmological constant. This may account for the missing factor of $1/4$ from their result. For completion, we can include other contributions of massless particles, namely the massless graviton, as well as $N_{1/2}$ massless spin-$1/2$ fermions, and $N_{1}$ spin-$1$ bosons, in addition to our $N=N_{0}$ spinless bosons. This gives
@@ 3.2.2 Long Distance Regime
[…]
(where $\Delta\mathcal{M}$ means the total amplitude minus the tree level result $i\kappa^{2}M_{1}^{2}M_{2}^{2}/(2q^{2})$). We then expand the left hand side in powers of $q$ around $q=0$, and on the right hand side we perform the integral, and then expand the result in powers of $q$. However, we must truncate this at some point. We choose to expand the left hand side out to 4 terms, i.e., $\sim q^{2},\,\sim q^{4},\,\sim q^{6}$, and $\sim q^{10}$, and we truncate the expansion on the right hand side at $b_{1}$. We then obtain 4 equations for 4 constants $A,a,p,b_{1}$. The value of $b_{1}$ is not expected to be accurate at this level of truncation. However, the value of $A,a,p$ is expected to be somewhat accurate. This truncated procedure yields the values: $A\approx 0.02798$, $a\approx 1.994$, and $p\approx 2.4499$. One could try working to higher order for more precision, however, then the integrals of the terms $b_{2}/(m_{i}^{2}r^{2})$ etc. don’t convert at small $r$. So this means this procedure cannot achieve full accuracy. A more technical treatment involves the method of steepest descent, which yields the slightly altered values of Frob:2016xte : $A=1/(16\sqrt{\pi})$, $a=2$, and $p=2.5$. This gives the potential to leading order in the large $mr$ regime as
@@
This is given as the dashed orange curve in Figure 1. We can compare these results to those of Burns:2014bva where we now disagree with their result (equation (V.57)) but do agree with Frob:2016xte .
@@ 3.3 Comparison to Solar System Bounds
Figure 2: Experimental bounds on the size of corrections to Newton’s potential $\alpha$ as a function of scale $\lambda$; taken from Ref. Adelberger:2003zx . Left side focuses on microscopic scales $10^{-6}m<\lambda<10^{-2}m$, while the right side focuses on macroscopic scales $10^{-2}m<\lambda<10^{14}m$.
Let’s compare this analysis to the massive correction given in (56) which was found in the limit that $mr\gg 1$. Although this form is not quite the same, since $\alpha$ in the above figure is taken as a constant, we can generalize this to (we set all masses equal here)
$\alpha\to\alpha(r)=\frac{NGm^{2}}{16\sqrt{\pi}(mr)^{3/2}}$ (58)
@@
In order to make a comparison we shall evaluate $\alpha$ at $r=\lambda$, which is the regime in which the observations are most sensitive. For example, for the lunar laser ranging (LLR), one can see in Figure 2 that the constraints are most sensitive at around $\lambda=$ few×$\times 10^{8}$ m, which is indeed the earth-moon distance. With this choice of $r=\lambda=1/(2m)$, then $\alpha$ is
@@
$\lambda\gtrsim 10^{4}\,km\,\,\,\,\,\,\mbox{for}\,\,\Lambda_{4}=1/(150\,km)$ (63)
@@
[…]
@@ 4 Conclusions
A possible extension of our work would be to include other types of massive fields within the loop and determine their contribution to Newton’s potential. Importantly, the $\alpha_{3}$ Wilson coefficient non-zero spin vanishes Bern:2021ppb . However, one can then use bounds on the quartic coupling, albeit involving logarthmic corrections.
@@ A.4 Fourier Transformations
$\int\frac{d^{3}q}{(2\pi)^{3}}\frac{e^{i\mathbf{q}\cdot\mathbf{r}}}{\absolutevalue{\mathbf{q}^{2}}}$ $=\frac{1}{4\pi r}$ (A14)
$\int\frac{d^{3}q}{(2\pi)^{3}}\frac{e^{i\mathbf{q}\cdot\mathbf{r}}}{\absolutevalue{\mathbf{q}}}$ $=\frac{1}{2\pi^{2}r^{2}}$ (A15)
$\int\frac{d^{3}q}{(2\pi)^{3}}e^{i\mathbf{q}\cdot\mathbf{r}}\ln(\mathbf{q}^{2})$ $=-\frac{1}{2\pi r^{3}}.$ (A16)
@@ Appendix B Effective action from the Amplitude
[…]

### v5

@@ 3.1.2 Cosmological Constant Counter Term
Let us first consider the cosmological constant, which appears in the action as $S\supset-\uprightintop\displaylimits d^{4}x\sqrt{-g}\,\Lambda_{0}$, where we have written the cosmological constant now as $\Lambda\to\Lambda_{0}$ to indicate that this is the bare value. Then consider the following expansion of the determinant of the metric
@@
$=\frac{i}{2}\left(\frac{-i\kappa}{2}\right)\mu^{4-d}\uprightintop\displaylimits\frac{d^{d}l}{(2\pi)^{d}}\frac{2l^{\mu}l^{\nu}-\eta^{\mu\nu}(l^{2}-m^{2})}{l^{2}-m^{2}}$
@@
Then using the Feynman rule that the linear term in the action gives $i\kappa\Lambda\eta^{\mu\nu}/2$, equation (44) therefore becomes
$\frac{i\kappa\Lambda_{R}}{2}\eta^{\mu\nu}=-\frac{i\kappa\Lambda_{0}}{2}\eta^{\mu\nu}+\frac{i\kappa m^{4}\eta^{\mu\nu}}{128\pi^{2}}\left(\frac{1}{\epsilon}+\log(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}})+\frac{3}{2}\right).$ (46)
@@ 3.2 Corrections to the Newtonian Potential
$V(\mathbf{r})=\frac{1}{2M_{1}}\frac{1}{2M_{2}}\uprightintop\displaylimits\frac{d^{3}q}{(2\pi)^{3}}\,e^{i\mathbf{q}\cdot\mathbf{r}}\mathcal{M}(\mathbf{q})$ (58)
@@
$V(r)=\frac{1}{2M_{1}}\frac{1}{2M_{2}}\frac{1}{2\pi^{2}}\displaylimits{}_{0}^{\infty}dq\,q^{2}\,\frac{\sin(q\,r)}{q\,r}\mathcal{M}(q).$ (59)
@@
Figure 1: Correction to the potential $\Delta V$ between point masses due to a scalar ($N=1$) running in a loop versus distance $r$. The black curve is the full numerical result from Fourier transforming Eq. (56). The dashed blue curve is the leading term of the small mrm\,r result given in Eq. (62). The dashed orange curve is the large mrm\,r result given in Eq. (66).
@@ 3.2.1 Short Distance Regime
If we take the massless limit we obtain just the single term $\Delta V=-NG^{2}M_{1}M_{2}/(20\pi r^{3})$; this is given as the dashed blue curve in Figure 1. This confirms Ref. Burns:2014bva , but is in disagreement with Ref. Faller:2007sy by a factor of $1/4$. We note that in Ref. Faller:2007sy there is no symmetry factor of $1/2$ in their bubble diagram, and they did not take into account our tadpole Feynman diagrams and the associated renormalization of the cosmological constant. This may account for the missing factor of $1/4$ from their result. For completion, we can include other contributions of massless particles, namely the massless graviton, as well as $N_{1/2}$ massless spin-$1/2$ fermions, and $N_{1}$ spin-11 bosons, in addition to our $N=N_{0}$ spinless bosons. This gives
@@ 3.2.2 Long Distance Regime
$\Delta\mathcal{M}(q)=2M_{1}\,2M_{2}\,4\pi\displaylimits{}_{0}^{\infty}dr\,r^{2}\,\frac{\sin(q\,r)}{q\,r}\,\Delta V(r)$ (65)
(where $\Delta\mathcal{M}$ means the total amplitude minus the tree level result $i\kappa^{2}M_{1}^{2}M_{2}^{2}/(2q^{2})$). We then expand the left hand side in powers of $q$ around $q=0$, and on the right hand side we perform the integral, and then expand the result in powers of $q$. However, we must truncate this at some point. We choose to expand the left hand side out to 4 terms, i.e., $\sim q^{2},\,\sim q^{4},\,\sim q^{6}$, and $\sim q^{10}$, and we truncate the expansion on the right hand side at $b_{1}$. We then obtain 4 equations for 4 constants $A,a,p,b_{1}$. The value of $b_{1}$ is not expected to be accurate at this level of truncation. However, the value of $A,a,p$ is expected to be somewhat accurate. This truncated procedure yields the values: $A\approx 0.02798$, $a\approx 1.994$, and $p\approx 2.4499$. One could try working to higher order for more precision, however, then the integrals of the terms $b_{2}/(m_{i}^{2}r^{2})$ etc. do not converge at small $r$. So this means this procedure cannot achieve full accuracy. A more technical treatment involves the method of steepest descent, which yields the slightly altered values of Frob:2016xte : $A=1/(16\sqrt{\pi})$, $a=2$, and $p=2.5$. This gives the potential to leading order in the large $mr$ regime as
@@
This is given as the dashed orange curve in Figure 1. We can compare these results to those of Burns:2014bva where we now disagree with their result (equation (V.57)) but do agree with Frob:2016xte  [^3] In Frob:2016xte , they corrected the large mass expansion found in Burns:2014bva which may have been due to the neglect of subleading terms in the expansion of special functions used (see section (4.3)).. However, numerically we do have agreement with Frob:2016xte in the massless and large mass limit.
We can also show direct equivalence between our results and theirs in the following manner. From the Fourier transformation of the amplitude to Newton’s potential, we can re-write it as (subtracting off Newton’s potential to focus on the correction)
$|\Delta V(r)=\frac{1}{16\pi^{2}M_{1}M_{2}}\frac{1}{ir}\displaylimits{}_{0}^{\infty}dq\>q(e^{iqr}-e^{-iqr})\mathcal{M}(q)e^{-\zeta\absolutevalue{q}}$ (67)
where we inserted a UV-regulator for completeness $|e^{-\zeta\absolutevalue{q}}$ to handle the divergences, but this will not be our primary focus nor impede our analysis and will drop it for further analysis. Notice that when we analytically continue the momentum $q\rightarrow ik$, our amplitude $\mathcal{M}(ik)$ will have a branch cut due to the logarithmic term between 2m<k<2m-2m<k<2m. Only considering positive mass terms, we can include these terms by computing the discontinuity across the branch cut namely
$\Delta V(r)=-\frac{1}{16\pi^{2}M_{1}M_{2}}\frac{1}{r}\displaylimits{}_{2m}^{\infty}dk\>ke^{-kr}\>\text{Disc}\>\mathcal{M}(ik)$ (68)
where the discontinuity is defined as $\text{Disc}\>F\equiv F(k+i0)-F(k-i0)$. Computing the discontinuity of $\mathcal{M}(ik)$ only involves the logarithm (which gives a factor of $\text{Disc}\log(\cdots)=2\pi i$ when the argument is less than 0, which occurs when $k>2m$). This gives us the following
$\text{Disc}\>\mathcal{M}(ik)=\sum_{i}\frac{\kappa^{4}M_{1}M_{2}}{3840\pi}\frac{\sqrt{k^{2}-4m_{i}^{2}}}{k^{5}}\left(28m_{i}^{4}-4m_{i}^{2}k^{2}+3k^{4}\right).$ (69)
Therefore, (68) takes the following form
$\Delta V(r)=-\sum_{i}\frac{\kappa^{4}M_{1}M_{2}}{61440\pi^{3}}\frac{1}{r}\displaylimits{}_{2m}^{\infty}dk\>e^{-kr}\sqrt{k^{2}-4m_{i}^{2}}\left(3-\frac{4m_{i}^{2}}{k^{2}}+\frac{28m_{i}^{4}}{k^{4}}\right).$ (70)
Here (70) matches that found in Burns:2014bva , and that of Frob:2016xte (their eq. (4.18) that was used to find the massless and massive limits).
@@ 3.3 Comparison to Solar System Bounds
Let’s compare this analysis to the massive correction given in (66) which was found in the limit that $mr\gg 1$. Although this form is not quite the same, since $\alpha$ in the above figure is taken as a constant, we can generalize this. However, to be precise, we need to specify a distribution for the masses, as we now explore.
3.3.1 All Masses Equal
Let us begin by setting all the masses equal $m=m_{\text{min}}=m_{1}=\ldots=m_{N}$. Then we can define an effective value for $\alpha$ from comparing to Eq. (66) as
$\alpha\to\alpha(r)=\frac{N\,G\,m^{1/2}}{16\sqrt{\pi}\,r^{3/2}}$ (72)
@@ 3.3.1 All Masses Equal
[…]
@@
Figure 2: Experimental bounds on the size of corrections to Newton’s potential $\alpha$ as a function of scale $\lambda$; taken from Ref. Adelberger:2003zx and Fischbach:1992fa . Left side focuses on microscopic scales $10^{-6}m<\lambda<10^{-2}m$, while the right side focuses on macroscopic scales $10^{-2}m<\lambda<10^{14}m$.
@@
$\lambda\gtrsim 10^{4}\,\mbox{km}\,\,\,\,\,\,\mbox{for}\,\,\Lambda_{4}=1/(150\,\mbox{km})$ (77)
@@
$\alpha_{\lambda}\simeq 8\times 10^{-5},\>\>$ $\text{at LIGO-Virgo scale}\>\lambda\sim 100\>\text{km}.$ (79)
@@
3.3.2 Distribution of Masses
Let us again consider the distributions of masses $\rho(m)$ that we introduced earlier in Eqs. (21) and (22). The corrected potential can be changed from the sum in Eq. (66) to
$\Delta V(r)=-\uprightintop\displaylimits dm\,\rho(m)\frac{G^{2}M_{1}M_{2}m^{3}}{16\sqrt{\pi}\,(mr)^{5/2}}e^{-2mr}$ (81)
By carrying out this integral for the power law and exponential distributions, taking the large $r$ limit, and identifying the effective $\alpha$, we obtain
$\alpha\to\alpha_{p}(r)={N\,G\over 32\sqrt{\pi}\,m_{\text{min}}^{1/2}\,r^{5/2}}\,\,\,\,\,(\mbox{Power Law})$ (82)
$\alpha\to\alpha_{e}(r)={N\,G\,m_{\text{min}}^{1/2}\over 32\sqrt{\pi}\,f\,r^{5/2}}\,\,\,\,\,(\mbox{Exponential})$ (83)
where again the scale $\lambda=1/(2m_{\text{min}})$ in both cases.
As above, we focus on the characteristic value relvant for observations at $r=\lambda=1/(2m_{\text{min}})$, where $\alpha$ takes the value
$\alpha_{\lambda,p}=\frac{N\,G\,m_{\text{min}}^{2}}{\sqrt{32\pi}}\,\,\,\,\,\mbox{(Power Law)}$ (84)
$\alpha_{\lambda,e}=\frac{N\,G\,m_{\text{min}}^{3}}{\sqrt{32\pi}\,f}\,\,\,\,\,\mbox{(Exponential)}$ (85)
We now update eq. (10) to a distribution as
${4\over\Lambda_{4}^{6}}=\uprightintop\displaylimits dm\,\rho(m)\,\frac{1}{(4\pi)^{2}}\frac{11}{75600}\frac{1}{m^{4}}\left(\frac{\kappa}{2}\right)^{2}$ (86)
(using $\alpha_{4}=4/\Lambda_{4}^{6}$). We carry out these integrals and solve for $N$ to obtain
$N={96768000\pi^{2}m_{\text{min}}^{4}\over 11\kappa^{2}\Lambda_{4}^{6}}\,\,\,\,\,\mbox{(Power Law)}$ (87)
$N={58060800\pi^{2}m_{\text{min}}^{3}\,f\over 11\kappa^{2}\Lambda_{4}^{6}}\,\,\,\,\,\mbox{(Exponential)}$ (88)
where in the exponential case, we have again taken the large $f\gg m_{\text{min}}$ limit, as this is the interesting regime. Inserting these expressions into $\alpha_{\lambda}$, and using the above reference values, we obtain an updated version of Eq. (76) as
$\alpha_{\lambda,p}=12000\left(\frac{m}{m_{0}}\right)^{6}\left(\frac{\Lambda_{4,0}}{\Lambda_{4}}\right)^{6}=12000\left(\frac{\lambda_{0}}{\lambda}\right)^{6}\left(\frac{\Lambda_{4,0}}{\Lambda_{4}}\right)^{6}\,\,\,\,\,\mbox{(Power Law)}$ (89)
$\alpha_{\lambda,e}=7300\left(\frac{m}{m_{0}}\right)^{6}\left(\frac{\Lambda_{4,0}}{\Lambda_{4}}\right)^{6}=7300\left(\frac{\lambda_{0}}{\lambda}\right)^{6}\left(\frac{\Lambda_{4,0}}{\Lambda_{4}}\right)^{6}\,\,\,\,\,\mbox{(Exponential)}$ (90)
Note that in the exponential case, the factor of $f$ has dropped out of this result (we have taken the $f\gg m_{\text{min}}$ limit). This may at first sight seem surprising, because in this limit the distribution becomes very broad and so one might expect a smaller value for $\alpha$ as it would seem to be dominated by heavier and heavier particles. However, note that we are fixing $\Lambda_{4}$ by altering $N$ accordingly. In Eq. (88) $N$ is taken to be proportional to $f$ to ensure there is no change in $\Lambda_{4}$. This results in the final $\alpha_{\lambda}$ being independent of $f$ in this limit.
We see that these values for $\alpha_{\lambda}$ are moderately larger than that for the delta-function distribution in Eq. (76). The pre-factor of 2500 has risen to 12000 and 7300, respectively, in each case. This makes such models all the more ruled out by solar system tests. However, we note that it is only a relatively modest change, since it is only an $\mathcal{O}(1)$ change in pre-factor. So the essential conclusions reached at the end of last subsection carry over.
3.4 Hawking Radiation
A concern in these models pertains to the huge number of degrees of freedom $N$ that one would need to integrate out to generate such large corrections. Recall the bound in Eq. (18), which indicates we may need to consider $N\sim 10^{82}$ or so (we took $\Lambda_{4}\sim 1/(8.6\,\mbox{km})$ for this estimate). With such a huge number of light particles, one should be concerned if black holes could even exist for a long period of time, rather than rapidly evaporating away via Hawking’s process. It is known that a black hole of mass $M$ has a lifetime from the production of $N$ massless (or nearly massless) particles of
$t_{BH}\approx{4\times 10^{67}\over N}\left(M\over M_{\odot}\right)^{3}\,\mbox{years}$ (91)
where $M_{\odot}$ is the solar mass. For moderate $N$ this is extremely long. But for $N\sim 10^{82}$ or so, this becomes a microscopic lifetime. Since we have observed black holes in mergers of masses $M=\mathcal{O}(10)M_{\odot}$, which must haved lived for a long time, this appears to be ruled out.
However, the particles we are studying are not massless. Let us consider the case that they all have the same mass $m$. Then, if we are in a regime in which the black hole temperature is $T_{BH}\lesssim m$, the output radiation is Boltzmann suppressed $\propto e^{-m/T_{BH}}$. The corresponding lifetime is enhanced from Eq. (91) by a factor $\sim e^{m/T_{BH}}$. Recall that the temperature of a black hole is related to its Schwarzschild radius as
$T_{BH}={1\over 4\pi R_{S}}={1\over 8\pi GM}.$ (92)
Now the lightest black holes observed (which have the highest temperature) have mass $M\approx 3\,M_{\odot}$. This corresponds to a temperature of
$T_{BH}\approx 1.7\times 10^{-12}\,\mbox{eV}\left(3\,M_{\odot}\over M\right).$ (93)
We note that this temperature is comparable to the kinds of masses we have been discussing. However, since it is of the same order, there is not much Boltzmann suppression.
On the other hand, as we increase precision and increase the value of the masses $m$ that we are sensitive to (which is better for a sensible EFT) then we have $T_{BH}\ll m$ and there is indeed appreciable suppression. By including the exponential enhancement in Eq. (91), and demanding that the black hole lifetime exceed $5\times 10^{9}$ years (as a useful reference), we find the masses would need to be bounded by
$m\gtrsim 1.7\times 10^{-12}\,\mbox{eV}\left(3\,M_{\odot}\over M\right)\ln\left[10^{-58}N\right]$ (94)
For $N\sim 10^{82}$, this corresponds to $m\gtrsim 9.6\times 10^{-11}$ eV. Consequently, to be consistent with causality ($\Lambda_{4}\geq 1.4\,m$) we need $\Lambda_{4}\geq 1.3\times 10^{-10}\,\mbox{eV}\approx 1/(1.5\,\mbox{km})$. This is yet another needed improvement in precision.
@@ 4 Conclusions
We also noted that to avoid over-production of Hawking radiation from black holes, one needs the mass of the scalars to be slightly heavier still. This requires slightly larger values for $\Lambda_{4}$, and so even further improvement in precision.
A possible extension of our work would be to include other types of massive fields within the loop and determine their contribution to Newton’s potential. Importantly, the $\alpha_{3}$ Wilson coefficient non-zero spin vanishes Bern:2021ppb . However, one can then use bounds on the quartic coupling, albeit involving logarithmic corrections.
@@ A.4 Fourier Transformations
$|\displaystyle\uprightintop\displaylimits\frac{d^{3}q}{(2\pi)^{3}}\frac{e^{i\mathbf{q}\cdot\mathbf{r}}}{\absolutevalue{\mathbf{q}^{2}}}$ $=\frac{1}{4\pi r}$ (A14)
$|\displaystyle\uprightintop\displaylimits\frac{d^{3}q}{(2\pi)^{3}}\frac{e^{i\mathbf{q}\cdot\mathbf{r}}}{\absolutevalue{\mathbf{q}}}$ $=\frac{1}{2\pi^{2}r^{2}}$ (A15)
$\uprightintop\displaylimits\frac{d^{3}q}{(2\pi)^{3}}e^{i\mathbf{q}\cdot\mathbf{r}}\ln(\mathbf{q}^{2})$ $=-\frac{1}{2\pi r^{3}}.$ (A16)
@@ Appendix B Effective action from the Amplitude
In this expansion, the non-local components of the Lagrangian (B11) cancel at first order in the expansion of (B12). This leads to terms on the order of $\Box/m^{2}$ that can become field re-definitions to the graviton propagator.